# MToP: A MATLAB Optimization Platform for Evolutionary Multitasking

### v4

@@ 2.3.2. Problem
All problems within MToP inherit from the problem base class <Problem.m>. The properties and methods of <Problem.m> are detailed in Table 7. As dimensions and upper and lower boundaries may vary across tasks, MToP offers a default unified search space mapping approach (Gupta et al., 2016). It’s important to note that while the unified search space approach serves as the default mapping method in MToP, alternative mapping techniques can also be implemented within specific algorithms.
@@
A simple problem implementation example is illustrated in Listing LABEL:list:prob. The maximum number of function evaluations, denoted as <maxFE>, for the problem can be specified within the class constructor. The function setTasks() is responsible for configuring the properties and evaluation function for each optimization task. Within this function, the number of tasks <T> is first defined, followed by a detailed setup for each task. Subsequently, parameters such as decision variable dimensions <D>, objective dimensions <M>, fitness function <Fnc>, lower bound <Lb>, and upper bound <Ub> are set individually for each task. For problems involving multiple tasks, <M> and <D> are represented as <vector> type, while <Fnc>, <Lb>, and <Ub> are represented as <cell> type. Notably, the function <Fnc> for each task takes the decision variable <Dec> as input and returns objective values <Obj> along with the constraint violation value <CV>. Moreover, for multi-objective optimization, an external function getOptimum() is required to define optimal solutions (i.e., reference points).
@@
With these standardized problem code patterns, all fully defined problems can be solved using the corresponding types of algorithms in MToP. The existing problems available in MToP are enumerated in Table 6. In addition to MTO problems, MToP also encompasses some single-task problems, the specifics of which can be found in the source code repository, but are not discussed herein.
Table 6. Benchmark and real-world optimization problems included in MToP. The table lists the problem name, case number, task type, dimension, optimization objective, task type, special case, and a brief description. SO and MO denote single-objective and multi-objective optimization, respectively. ST, MT and MaT represent single-task, multi-task and many-task, respectively.
Problem Case Task Dimension Objective Task Special Description
CEC17-MTSO (Da et al., 2017) 9 2 50 SO MT Complete/partial/no global optima interaction and high/medium/low inter-task similarity
CEC19-MaTSO (Feng et al., 2019a) 6 2-50 50 SO MaT Single-objective many-task optimization problems
WCCI20-MTSO (Feng et al., 2020) 10 2 50 SO MT Complex single-objective MTO problems
WCCI20-MaTSO (Feng et al., 2020) 10 2-50 50 SO MaT Complex single-objective many-task optimization problems
CMT (Li et al., 2022a) 9 2 Any SO MT Constrained Large/partial/no feasible regions interaction and different inter-task similarity
C2TOP&C4TOP (Li et al., 2022c) 51 2/4 50 SO MT Competitive Different optimal objective values among competitive tasks
C-CPLX (Li et al., 2023a) 20 2 50 SO MT Competitive Complex competitive MTO problems
LSMaTSO (Li et al., 2024b) 5 2-50 300 SO MaT Large-scale many-task optimization problems
CEC17-MTMO (Yuan et al., 2017) 9 2 50 MO MT Complete/partial/no global optima interaction and high/medium/low inter-task similarity
CEC19-MTMO (Feng et al., 2019a) 10 2 50 MO MT Complex multi-objective MTO problems
CEC19-MaTMO (Feng et al., 2019a) 10 2-50 50 MO MaT Multi-objective many-task optimization problems
WCCI20-MaTMO (Feng et al., 2020) 10 2-50 50 MO MaT Complex multi-objective many-task optimization problems
CEC21-MTMO (Feng et al., 2021b) 10 2 50 MO MT More complex multi-objective multi optimization problems
CMOMT (Li et al., 2024g) 12 2/3 30-60 MO MT Competitive multiobjective MTO problems
Synthetic Functions 9 1 Any SO ST Ackley, Elliptic, Griewank, Rastrigin, Rosenbrock Schwefel, Sphere, Weierstrass
CEC06-CSO (Liang et al., 2006) 24 1 2-24 SO ST Constrained CEC 2006 competition on constrained single-objective optimization
CEC10-CSO (Mallipeddi and Suganthan, 2010a) 18 1 10/30 SO ST Constrained CEC 2010 competition on constrained single-objective optimization
CEC17-CSO (Wu et al., 2017) 28 1 2-100 SO ST Constrained CEC 2017 competition on constrained single-objective optimization
CEC17-SO (Awad et al., 2017) 29 1 2-100 SO ST CEC 2017 competition on bound constrained single-objective optimization
CEC20-RWCO (Kumar et al., 2020) 57 1 2-158 SO ST Constrained CEC 2020 competition on real-world constrained single-objective optimization
CEC22-SO (Kumar et al., 2022) 12 1 10/20 SO ST CEC 2022 competition on bound constrained single-objective optimization
PEPVM (Li et al., 2023b) 1 3 7 SO MT Parameter extraction of photovoltaic models as MTO problem
SCP (Li et al., 2022c) 1 Any 60-120 SO MT Competitive Sensor coverage problem as competitive MTO problem
OPF (Li et al., 2024c) 5 2 24/33 SO MT Constrained Optimal power flow as constrained MTO problem
PKACP (Jiang et al., 2023a) 1 Any Any SO MaT Planar kinematic arm control problem as many-task optimization problem
MO-SCP (Li et al., 2023b) 2 Any 60-120 MO MT Competitive Multiobjective sensor coverage problem as competitive multiobjective MTO problem
SOPM (Li and Gong, 2025) 2 3 30 MO MT Constrained Synchronous optimal pulse-width modulation as constrained multiobjective MTO problem
MO-OPF (Li and Gong, 2025) 4 2 24/33 MO MT Constrained Multiobjective optimal power flow as constrained multiobjective MTO problem
OPF-CMTMO (Li et al., 2024g) 4 2 11/24 MO MT Competitive Competitive multitask optimal power flow with thermal and wind-solar power
@@
N Population size for each task
@@ 11 Prob = Prob@Problem(varargin);
30function optimum = getOptimum(Prob)
@@ 2.3.3. Performance Metric
11% Data for shown in the GUI table
12result.RowName = {MTOData.Problems.Name};
13result.ColumnName = {MTOData.Algorithms.Name};
14result.TableData = CalculateTableData(MTOData);
@@
16% Data for shown in the GUI convergence plot
17result.ConvergeData = CalculateConvergeData(MTOData);
@@
19% Data for shown in the GUI Pareto plot
20result.ParetoData = CalculateParetoData(MTOData);
@@
Unlike the implementation of algorithm and problem classes, metric codes in MToP are defined as functions. An illustrative example of metric implementation is presented in Listing LABEL:list:metric. The input parameter of the function is <MTOData>, which is generated during experimental execution. The function returns <result>, comprising properties such as <Metric>, <RowName>, <ColumnName>, <TableData>, <ConvergeData>, and <ParetoData>. The <Metric> property can take values of either ’Min’ or ’Max’, indicating whether a smaller or larger metric value is preferable. <RowName>, <ColumnName>, and <TableData> are utilized to present metric results in the GUI table. On the other hand, <ConvergeData> and <ParetoData> are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively. Note that specific metric calculation functions are not elaborated here for the sake of simplicity. The metrics currently integrated into MToP are enumerated in Table 8.
Table 8. Metrics included in MToP. The table lists the metric name, optimization objective, task type, and a brief description. SO and MO denote single-objective and multi-objective optimization, respectively. ST and MT represent single-task and multi-task, respectively
Metric Objective Task Description
Obj SO ST/MT Objective value for each task
Obj (AV) SO MT Average Obj for all tasks
Obj (UV) SO MT Unified average Obj for all tasks
Obj (MTS) SO MT Multitask score of Obj for all tasks
Obj (CMT) SO MT Competitive multitask Obj for all tasks
Obj (NBR) SO MT Number of best Obj result for all tasks
CV SO ST/MT Constraint violation for each task
FR SO ST/MT Feasible rate for each task
HV MO ST/MT Hypervolume for each task
HV (MTS) MO MT Multitask score of HV for all tasks
HV (CMT) MO MT Competitive multitask HV for all tasks
IGD MO ST/MT Inverted generational distance for each task
IGD (AV) MO MT Average IGD for all tasks
IGD (MTS) MO MT Multitask score of IGD for all tasks
IGD (CMT) MO MT Competitive multitask IGD for all tasks
IGD+ MO ST/MT Improved IGD plus for each task
IGD+ (MTS) MO MT Multitask score of IGD+ for all tasks
IGD+ (CMT) MO MT Competitive multitask IGD+ for all tasks
Spread MO ST/MT Spread metric for each task
Spread (CMT) MO MT Competitive multitask spread for all tasks
Run Time SO/MO ST/MT Algorithm running time for all tasks
3. Performing Experiments with MToP
@@
To launch MToP, start by running the script file <mto.m> located in the root directory. This action will initialize the MToP GUI interface, which is illustrated in Fig. 6. The GUI interface of MToP requires MATLAB R2020b or later versions to run, while command-line execution can be done with any version. In addition, if parallel execution is required, MATLAB Parallel Computing Toolbox is needed.
@@ 3.1. Testing Algorithms and Problems
As depicted in Fig. 6 (a), the test module interface is structured into left, center, and right sections. The left section facilitates the selection and configuration of algorithms and problems. Upon choosing the desired labels such as <Task>, <Objective>, and <Special> from the drop-down buttons, the options available for <Algorithm> and <Problem> are updated accordingly. Selecting specific algorithms and problems reveals their parameters in the edit box, which can be edited by clicking on them. Once parameters are adjusted, the new values are applied to the algorithm and problem objects through the setParameter() method. Algorithm parameters typically include evolutionary operator settings and other customizable parameters, while problem parameters usually consist of population size <N>, maximum function evaluations <maxFE>, and other specific parameters.
In the center section of the MToP GUI, users can execute the test experiment and visualize the results graphically. Upon completion of the algorithm execution on the chosen problem, the metric results of the problem are displayed in the right section of the test module.
3.2. Executing Comparative Experiments
@@
The GUI of the experiment module is displayed in Fig. 6 (b), featuring a three-column layout. The left and middle columns comprise the experimental setup, while the right column displays the results.
In the first column, users can configure the basic experiment settings. The <No. of Runs> textbox specifies the number of independent repeat runs for each algorithm on each problem. The <No. of Results> textbox determines the number of results saved during the evolutionary process of each algorithm, sampled uniformly according to the number of generations. The <Parallel> button toggles parallel processing, allowing independent experiments to run simultaneously across multiple process. The number of concurrent experiments depends on the number of processor cores. If the number of parallel experiments exceeds the available cores, MATLAB automatically handles the scheduling. For multiple algorithms and problems, each algorithm processes a problem in parallel. After one algorithm finishes, the next starts, repeating until all algorithms have processed the problem, then moving to the next problem. Hence, the number of algorithms and problems is not inherently limited. The drop-down buttons <Task>, <Objective>, and <Special> filter algorithms and problems based on predefined labels, similar to the test module. Users can add selected algorithms and problems by clicking the <Add> button after choosing them from the respective drop-down boxes at the bottom.
In the second column, the names and parameters of the added algorithms and problems can be edited by clicking on them. Once all configurations are set, clicking the green <Start> button initiates the experiment execution. Users can also pause or stop the experiment using the <Pause> and <Stop> buttons if needed. Once the experiment is completed, the table’s contents become available in the results area on the right part of the experiment module. Users can read and save the original data of the experiment using the <Load Data> and <Save Data> buttons.
In the results area, the first function row contains buttons for operating on the experimental data. The first button, labeled <Save Table>, exports the table results to xlsx, csv, and tex files. The subsequent buttons labeled <Converge> control the convergence y-lim types and generate convergence plots based on the selected results from the table. The <Pareto> button draws the Pareto front and non-dominated solutions based on the selected results.
The edit box controls the formatting type of the table result. Following this, the drop-down box allows users to select performance metrics, and subsequent drop-down boxes determine the result display type, statistical test method, base algorithm, and result highlighting type, respectively. After selecting the statistical test method, calculated statistical results denoted by “+ / - / =” followed by each metric result can also be displayed in the table. MToP offers two non-parametric methods, the Wilcoxon signed rank test and the Wilcoxon rank-sum test, for statistical tests. The base algorithm drop-down button controls the base algorithm in the statistical test, while the highlight button highlights the best and worst algorithm’s metric results for each problem or task.

### v5

@@ 3.3.2. Problem
All problems within MToP inherit from the problem base class Problem.m. The properties and methods of Problem.m are detailed in Table 4. As dimensions and upper and lower boundaries may vary across tasks, MToP offers a default unified search space mapping approach as Eq. (2). It’s important to note that while the unified search space approach serves as the default mapping method in MToP, alternative mapping techniques can also be implemented within specific algorithms.
@@
A simple problem implementation example is illustrated in Listing LABEL:list:prob. The maximum number of function evaluations, denoted as maxFE, for the problem can be specified within the class constructor. The function setTasks() is responsible for configuring the properties and evaluation function for each optimization task. Within this function, the number of tasks T is first defined, followed by a detailed setup for each task. Subsequently, parameters such as decision variable dimensions D, objective dimensions M, fitness function Fnc, lower bound Lb, and upper bound Ub are set individually for each task. For problems involving multiple tasks, M and D are represented as vector type, while Fnc, Lb, and Ub are represented as cell type.
Notably, the function Fnc for each task takes the decision variable Dec as input and returns objective values Obj along with the constraint value Con. The problem’s base class evaluate() method manages this process by using the specific function handle defined in Fnc to perform the calculation for the requested task.
Moreover, for multi-objective optimization, the external function getOptimum() is provided. It defines the true optimal solutions (i.e., true Pareto front) if known, which is required for metrics like IGD. For problems where the true front is unknown (e.g., real-world applications), this function is instead used to provide the reference point for calculating metrics such as hypervolume (HV).
@@
With these standardized problem code patterns, all fully defined problems can be solved using the corresponding types of algorithms in MToP. The existing problems available in MToP are enumerated in Table LABEL:tab:prob-mt of the supplementary files.
@@
N Default population size for each task
@@ 11 Prob = Prob@Problem(varargin);
30function optimum = getOptimum(Prob) % optional
@@ 3.3.3. Performance Metric
11 % Data for shown in the GUI table
12 result.RowName = {MTOData.Problems.Name};
13 result.ColumnName = {MTOData.Algorithms.Name};
14 result.TableData = CalculateTableData(MTOData);
@@
16 % Data for shown in the GUI convergence plot (optional)
17 result.ConvergeData = CalculateConvergeData(MTOData);
@@
19 % Data for shown in the GUI Pareto plot (optional)
20 result.ParetoData = CalculateParetoData(MTOData);
@@
Unlike the implementation of algorithm and problem classes, metric codes in MToP are defined as functions. An illustrative example of metric implementation is presented in Listing LABEL:list:metric. The input parameter of the function is MTOData, which is generated during experimental execution. The function returns result, comprising properties such as Metric, RowName, ColumnName, TableData, ConvergeData, and ParetoData. The Metric property can take values of either Min or Max, indicating whether a smaller or larger metric value is preferable. RowName, ColumnName, and TableData are utilized to present metric results in the GUI table. On the other hand, ConvergeData and ParetoData are employed to exhibit metric convergence results for convergence plots and non-dominated solutions for Pareto plots, respectively. Note that specific metric calculation functions are not elaborated here for the sake of simplicity. The metrics currently integrated into MToP are enumerated in Table LABEL:tab:metric of the supplementary files.
4. Guidelines for Using MToP
@@
To launch MToP, start by running the script file mto.m located in the root directory. This action will initialize the MToP GUI interface, which is illustrated in Fig. 6. The GUI interface of MToP requires MATLAB R2022b or later versions to run, while command-line execution can be done with any version.
@@ 4.1. Testing Algorithms and Problems
As depicted in Fig. 6 (a), the Test Module interface is structured into left, center, and right sections. The left section is dedicated to configuration, the center section serves as the main display for visualizations, and the right section shows the final metric results after a run. This module is designed for preliminary analysis, debugging, and qualitative visualization of a single algorithm’s performance on a single problem.
4.1.1. Viewing Problem Characteristics
Figure 7. Problem characteristics visualization in the Test Module of MToP GUI.
Before executing an algorithm, researchers can use the Test Module to visualize the characteristics of the selected problem. As shown in Fig. 7, after selecting a problem (e.g., CEC17-MTMO1-CI-HS), the central display area provides a drop-down menu (initially labeled Tasks Figure (1D Unified) in the figure). Clicking this menu reveals various visualization options. These options vary by problem type but include plotting 1D or 2D function landscapes to understand the search space, visualizing feasible regions for constrained problems, and displaying the true Pareto front for multi-objective tasks after running the algorithm. This allows for an initial assessment of the problem’s difficulty and features prior to optimization.
4.1.2. Setting Algorithm and Problem Parameters
Figure 8. Algorithm and problem parameter setting in the Test Module of MToP GUI.
The left section of the module handles the selection and configuration of algorithms and problems. As illustrated in Fig. 8, users first select the desired algorithm (e.g., MO-MFEA) and problem (e.g., CEC17-MTMO1-CI-HS) from the respective lists. Filtering options (e.g., Task, Objective, Special) help narrow down the choices. Once selected, their default parameters appear in the text boxes below. To modify a parameter, the user can double-click the parameter’s name in the list, which opens an editable field (as shown in the figure for Change Name and Change Parameter). For example, the algorithm’s name can be changed, or a specific numeric parameter like the problem’s population size N can be adjusted. These changes are applied to the objects when the experiment is started by pressing the Start button in the central panel.
4.1.3. Visualizing Algorithm Behavior
After configuring the parameters and starting the run, MToP offers dynamic display utilities such as Draw Obj and Draw Dec to enable researchers to explore algorithm behavior in real-time. Figure 9 illustrates an example of population variation of MO-MFEA on CEC17-MTMO4, and compares its final state with that of MTDE-MKTA, showcasing the evolution in both objective space and decision space. CEC17-MTMO4 is a multi-objective problem with two tasks, each comprising two objectives. The optimal Pareto set of the first task contains diversity dimension $1$ with optimal value range $[0,1]$ and convergence dimension $[2,50]$ with optimal value $0.5$. The optimal Pareto set of the second task contains diversity dimension $1$ with optimal value range $[0,1]$, convergence dimension $[2,40]$ with optimal value $0.5$, and convergence dimension $[41,50]$ with optimal value $0.5005$. Through these dynamic visualizations, researchers can intuitively grasp the algorithm’s operational behavior and delve deeper into its search dynamics.
(a) Gen=1
(b) Gen=200
(c) Gen=500
(d) Gen=700
(e) Gen=1000
(f) MTDE-MKTA
(g) Gen=1
(h) Gen=200
(i) Gen=500
(j) Gen=700
(k) Gen=1000
(l) MTDE-MKTA
Figure 9. Schematic of population snapshots on CEC17-MTMO4 across generations. The top row (a)-(f) shows the objective space, where the blue dots are the algorithm’s non-dominated solutions and the black line is the true Pareto Frontier (PF). The bottom row (g)-(l) shows the decision space, where each black polyline represents the 50-dimensional decision vector of one individual in the population. Panels (a)-(e) and (g)-(k) depict MO-MFEA at Gen=1, 200, 500, 700, and 1000. Panels (f) and (l) show the example of MTDE-MKTA at Gen=1000.
In this example, during the initial phase of evolution (Gen=1), the population initialization is dispersed across the decision space and poorly situated in the objective space. As evolution progresses through the first and middle phases (Gen=200), individual gene knowledge transfer leads to populations swiftly converging to more favorable positions in the objective space, albeit at the expense of reduced diversity. Subsequently, in the middle and late stages (Gen=500 to 700), the population gradually emphasizes diversity. By Gen=1000, MO-MFEA achieves the optimal Pareto frontier on the first task but becomes trapped in local optima on the second task. The schematics provide the key insight: the population for the second task is clustered in the similar decision space region as the population for the first task. However, the true optimal regions of these two tasks do not precisely overlap in the decision space. This observation strongly suggests the presence of negative knowledge transfer: the strong convergence of Task 1 has incorrectly pulled the Task 2 population into its own optimal region, causing it to converge on a local optimum for Task 2. This visualization functionality in MToP is crucial for diagnosing such algorithmic failures.
Furthermore, MToP can be used to validate the effectiveness of remedial approaches. For instance, Figure 9 (f) and (l) show the final population snapshot of MTDE-MKTA, an algorithm with multiple knowledge transfer mechanisms designed to mitigate this issue. In contrast to MO-MFEA, MTDE-MKTA successfully converges to the true optimal Pareto frontier for both tasks. This is further confirmed in the decision space, where the inset plot clearly shows the convergence to the correct optimal values ($0.5$ for dimensions $[2,40]$ and $0.5005$ for $[41,50]$), thus overcoming the negative transfer problem. In MToP, the decision space population function applies not only to multi-objective problems but also to single-objective problems.
4.2. Performing Comparative Experiments
@@
The GUI of the Experiment Module, shown in Fig. 6 (b), features a three-column layout: the left and middle columns are for experimental setup, and the right column is for results analysis. The detailed steps for conducting an experiment are illustrated in Fig. 10.
Figure 10. Experiment execution steps in the Experiment Module of MToP GUI.
4.2.1. Setting Experiment Parameters
Experiment parameters are set in the Set Exp Parameter panel (Step 1). This includes setting the number of Repetitions (independent runs), the Rand Seed for reproducibility, and the Data Length (the number of convergence points to save per run). Toggles for Parallel execution and SaveDec (to save decision variables) are also located here. Enabling Parallel allows independent runs to execute simultaneously. The number of concurrent processes depends on the system’s processor cores, and MATLAB automatically handles the scheduling. The typical execution flow processes one problem at a time, running all added algorithms on it in parallel before moving to the next problem.
4.2.2. Selecting Algorithms and Problems
As shown in Fig. 10 (Steps 2-4), the user first selects the algorithms and problems to be included in the experiment. To simplify this process, users can use the tag filters (Step 2) such as Task, Objective, and Constrained to narrow down the lists based on predefined labels. From the Algorithm List (Step 3), the user selects one or more algorithms and clicks the Add button to move them to the Added Algorithm Tree. An identical process is used to select problems from the Problem List and add them to the Added Problem Tree (Step 4).
As shown in Steps 3 and 4, users can set parameters for specific algorithms and problems. By clicking on an item in the Added Algorithm Tree or Added Problem Tree, its individual parameters (e.g., population size N, operator parameters RMP) are displayed in the text box below, where they can be directly edited.
To assist researchers in this selection, especially when starting a new study, Table 5 provides a curated guide to benchmark suites and high-performing algorithms. The listed benchmarks are widely used and recognized within the EMT community, serving as standard testbeds. The recommended algorithms are not arbitrary; they represent the top 5 algorithms based on the Friedman ranking of their average metric performance over 30 independent runs on each benchmark set. The full experimental validation for this ranking is detailed later in Section 5.1. This table serves as an excellent starting point for researchers and can also offer valuable insights when selecting algorithms for new real-world problems. Comprehensive lists, references, and detailed descriptions of all problems, algorithms, and metrics available in MToP are provided in Table LABEL:tab:prob-mt, Table LABEL:tab:algo-mt, and Table LABEL:tab:metric, respectively, in the supplementary files.
Table 5. Recommended benchmarks, algorithms and metrics for different MTO problem categories in MToP. The selected algorithms are based on their Friedman rankings on each benchmark. Detailed references and descriptions of these algorithms and benchmarks can be found in supplementary files.
Benchmark Recommended Algorithms (selected by Friedman rankings) Recommended Metric(s)
Single-objective Multi-task
CEC17-MTSO MFEA-GHS, MTES-KG, MFMP, MTDE-ADKT, MTEA-HKTS Obj
WCCI20-MTSO MTES-KG, MFMP, MTDE-ADKT, MTEA-PAE, MFEA-DGD Obj
Constrained Single-objective Multi-task
CMT MTEA-PAE, MTES-KG, CEDA-MP, MFEA-GHS, MTEA-AD Obj, CV
Competitive Single-objective Multi-task
C2TOP&C4TOP MTSRA, MFMP, MTES-KG, MTEA-HKTS, DEORA Obj (CMT)
C-CPLX MTES-KG, MFMP, MTSRA, MTEA-PAE, MTEA-HKTS Obj (CMT)
Single-objective Many-task
CEC19-MaTSO DTSKT, TNG-SNES, MTES-KG, MTEA-HKTS, SBCMAES Obj (MTS), Obj (UV)
WCCI20-MaTSO DTSKT, TNG-SNES, MTES-KG, SBCMAES, EMaTO-MKT Obj (MTS), Obj (UV)
Multi-objective Multi-task
CEC17-MTMO MO-MTEA-PAE, MTDE-MKTA, MTEA-DCK, MTEA-D-DN, KR-MTEA IGD+, HV
CEC19-MTMO MTDE-MKTA, MTEA-DCK, MTEA-D-TSD, MO-MTEA-PAE, MTEA-D-DN IGD+, HV
CEC21-MTMO MTEA-D-TSD, MO-MTEA-PAE, MTEA-DCK, MO-MTEA-SaO, MTDE-MKTA IGD+, HV
Competitive Multi-objective Multi-task
CMOMT RVC-MTEA, MO-MTEA-PAE, MO-MTEA-SaO, MO-EMEA, MO-MFEA-II IGD+ (CMT), HV (CMT)
4.2.3. Running Experiment and Calculating Metrics
Once all configurations are set, the experiment is initiated by clicking the Start button in the Run / Load Experiment panel (Step 5). The Pause and Stop buttons can be used to control the execution. After the experiment completes, the user can move to the results panel (Step 6) to analyze the performance. The first step is to select a Metric (e.g., Obj) and a Show Type (e.g., Mean&Std). MToP then calculates the results and populates the main table.
Table 5 also provides guidance on which metrics are most appropriate, as the choice is critical for correct analysis:
  • For standard multi-task problems (e.g., CEC17-MTSO, CEC19-MTMO), the number of tasks is small, making it practical and recommended to analyze per-task metrics directly, such as Obj, IGD+, or HV. For multi-objective metrics, IGD+ is generally recommended over the traditional IGD as it provides a more comprehensive and stable measure of both convergence and distribution. Furthermore, the HV metric is also provided, as it is a valuable indicator that does not require a known true Pareto front, making it suitable for real-world problems where the optimum is unknown.
  • For constrained multi-task problems (e.g., CMT), the evaluation must account for both feasibility and objective quality. The standard approach is to report the Obj (objective value) for runs that find at least one feasible solution, but to report the CV (constraint violation) for runs that fail to find any feasible solutions.
  • For competitive multi-task problems (e.g., C-CPLX, CMOMT), per-task analysis is not correct, and aggregated metrics like Obj (CMT) or IGD+ (CMT) are necessary to capture the overall trade-off across tasks.
  • For many-task problems (e.g., CEC19-MaTSO), the large number of tasks makes reporting per-task metrics impractical. Here, we recommend using widely-adopted summary metrics such as the multi-task score Obj (MTS) (relative performance across tasks) or the unified value Obj (UV) (adjusted absolute performance across tasks).
  • For all problem types, the Run Time metric can be used to compare the computational efficiency and execution time of different algorithms.
4.2.4. Performing Statistical Tests
The results panel (Step 6) also includes a suite for statistical analysis. MToP provides two distinct non-parametric approaches for performance comparison: direct pairwise tests and a global omnibus test with post-hoc analysis.
First, for direct pairwise comparisons, MToP offers the Wilcoxon rank-sum test and the Wilcoxon signed-rank test. When a user selects one of these methods, MToP performs a direct statistical comparison between each algorithm and a user-selected Base Algorithm. MToP’s logic then annotates the results table with “+” (significantly better), “-” (significantly worse), or “=” (no significant difference) based on a $p$-value threshold of $0.05$. It is important to note that these are uncorrected pairwise tests intended for exploratory analysis.
Second, for a more robust global comparison, MToP provides the Friedman test. This is an omnibus test that first checks for any statistically significant differences among the entire set of selected algorithms. Users can choose to run this test on the mean results across all runs (mean) or on the complete dataset combining all runs (all reps). Upon selection, MToP automatically performs the global Friedman test and then immediately conducts a post-hoc test. This post-hoc analysis calculates the $z$-value and corresponding $p$-value to compare the mean rank of each algorithm against the mean rank of the Base Algorithm. The results table is then populated with the mean rank for each algorithm and the $p$-value for its comparison against the base, allowing for a statistically grounded, rank-based assessment.
It is a known characteristic of null-hypothesis significance testing that with a sufficiently large number of replications, even trivial performance differences can be deemed statistically significant. This does not necessarily imply practical significance. MToP is therefore designed to facilitate a multi-faceted analysis. MToP provides both relative performance comparisons (via the Wilcoxon and Friedman statistical test results) and crucial absolute performance measures (the Mean&Std values). Researchers are thus equipped to interpret the statistical results in conjunction with the absolute performance data and the MToP’s visualizations (e.g., Converge plots) to make a more informed judgment about the practical significance of the observed differences.
4.2.5. Exporting Experiment Data and Results
MToP provides two primary methods for saving data, located in the Run / Load Experiment panel (Fig. 10, Step 5). The Save Data and Load Data buttons are used to save or reload the entire experimental session (the raw MTOData object), which allows for experiments to be reloaded and analyzed later.
For saving processed results, the Export button provides more granular control. Clicking this button opens a new dialog window offering three distinct export options:
  • Current Table (tex, xlsx, csv): This option saves the data exactly as it is currently displayed in the main results table, including the calculated metrics and any statistical annotations (e.g., “+ / - / =” and “Ranking”). It supports tex, xlsx, and csv formats for easy integration into publications and spreadsheets.
  • IOHanalyzer Data (csv): This option exports the complete convergence history (not just the final values) for all selected runs into a csv format. This format is specifically structured for use with the IOHanalyzer (Wang2022IOHanalyzer), facilitating advanced external analysis.
  • Best Decision Variable (mat): This option extracts the decision variables of the best-found solutions of all algorithms and repetitions and saves them to a .mat file for further inspection or use in other applications. For single-objective MTO problems, the best solution is determined by the lowest objective value. For multi-objective MTO problems, the best solutions are a set of Pareto sets (Liu2024Finding; Choong2022Jack).
4.2.6. Visualizing Metric Convergence and Pareto Front
Finally, the module offers integrated plotting tools at the bottom of the results panel (Step 6). By selecting one or more rows and columns in the results table, the user can generate plots. Clicking the Converge button generates a convergence plot for the selected metric results, with a Log toggle available to switch the y-axis to a logarithmic scale, and a Range toggle to display the 95% confidence interval around the mean curve. For multi-objective problems, clicking the Pareto button will plot the final acquired non-dominated solutions (Pareto fronts). Additionally, the convergence data can be exported in a format compatible with IOHanalyzer (Wang2022IOHanalyzer) for further external analysis and visualization.
4.3. Processing Experiment Data
MToP provides a dedicated Data Process Module within the GUI for merging and splitting experimental data files. This module allows users to combine multiple MTOData.mat files into a single comprehensive dataset or to partition a large dataset into smaller, more manageable segments based on specific criteria such as algorithms, problems, or repetitions. The merging function is particularly useful for aggregating results from separate experimental runs, while the splitting function aids in isolating subsets of data for focused analysis. Users can access this module directly from the main GUI, where they can select the desired operation (merge or split), specify input files, and define output parameters. The processed data is saved in the standard MTOData.mat format, ensuring compatibility with other modules within MToP.
Figure 11. Data processing steps in the Data Process Module of MToP GUI.
4.3.1. Loading and Viewing Data Files
As shown in the Load / Save Data panel of Fig. 11, the user begins by clicking the Load Data button. This opens a file dialog to select one or more MTOData.mat files, which are then added to the Data Tree. By clicking on a loaded data object, its high-level structure (e.g., Reps, Algorithms, Problems) is displayed. Users can further click on these fields, such as Problems, to view the detailed contents in a table, allowing for verification of the data before processing. The Delete Data button can also be used to remove any unwanted or intermediate data objects from the Data Tree.
4.3.2. Splitting Experiment Data
The Split Data panel in Fig. 11 illustrates the partitioning functionality. To divide a large dataset, the user selects a data object from the Data Tree. MToP provides three splitting options: Reps Split, Algorithm Split, and Problem Split. For instance, selecting Algorithm Split will parse the dataset and create new, separate data objects for each algorithm it contains (e.g., MTOData (Split Algorithm: MFEA)). These new objects then appear in the Data Tree, ready for isolated analysis or saving.
4.3.3. Merging Experiment Data
Conversely, the Merge Data panel demonstrates how to combine datasets. The user selects two or more data objects from the Data Tree that share compatible settings. For example, to aggregate independent runs, the user can select multiple files and click Reps Merge. This creates a single new dataset (e.g., Data (Merge Reps)) with a larger total number of repetitions. The Algorithm Merge and Problem Merge buttons function similarly, allowing for the consolidation of data across algorithms or problems. Note that when merging, MToP checks for consistency in experiment settings to ensure valid combinations (e.g., same problems and reps when merging algorithms).
4.3.4. Adjusting Data Precision
The module also provides tools for changing data precision. Users can select a data object and specify the desired number of decimal places for metric values. This is particularly useful for reducing file size or standardizing data formats before saving.
4.3.5. Saving Processed Data
After all processing operations (splitting, merging, or reordering) are complete, the new or modified data objects can be exported. As shown in the Load / Save Data panel, the user selects the desired data object from the Data Tree (e.g., Data (Merge Reps)) and clicks the Save Data button to save it as a new MTOData.mat file.